Grade 11Algebra

what is real valued function? plzz..xplain with a simple example...!!

Ranjita yadav
14 Years agoGrade 11

2 Answers

Ashwin Muralidharan IIT Madras
ApprovedApproved Tutor Answer14 Years ago

Hi Ranjita,

 

Real Valued functions are those functions that have real arguments (domain) and return a real value upon evaluation (range).

y = f(x).

x is real and y is real.

There could be numerous examples.

y=x^2, y=sinx.... etc.

 

Regards,

Ashwin (IIT Madras).

Ranjita yadav
14 Years ago

so how other functions are differnt from this..?? real domain & range can also be in other functions too..!!